OpenStudy (anonymous):

The radius of the earth is 6370Km and the radius of mars is 3440Km. If an object weights 200N on earth, what will be the gravity on mars? Mars has a mass of 0.11 times the mass of the earth.

OpenStudy (anonymous):

mg =200 g(mars) = GM(mars)/R(mars)^2 W(mars) = mg(mars)

OpenStudy (anonymous):

how do you get there this expression? g(mars) = GM(mars)/R(mars)^2

OpenStudy (anonymous):

g of any planet is given by this formula

OpenStudy (anonymous):

and G is a constant, isnt it?

OpenStudy (anonymous):

if the objects mass is m and planets mass is M and radius R the force bw them is GMm/R^2 so the acceleration due to gravity on the planet is F/m = GM/R^2

OpenStudy (anonymous):

G is the Universal Gravitational Constant

OpenStudy (anonymous):

F=GMm/R^2

OpenStudy (anonymous):

yes....now F/m is acceleration

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so now ur doing gravitation? how old r u?

OpenStudy (anonymous):

Im helping my newphew. He is finishin high school.

OpenStudy (anonymous):

im 22 year old.

OpenStudy (anonymous):

and what r u studying?

OpenStudy (anonymous):

Electronics and telecomunications. At the university. Im working on my thesis proyect. I received physics when I was in 2d semester.

OpenStudy (anonymous):

thesis on what topic?

OpenStudy (anonymous):

Pattern recognition :)

OpenStudy (anonymous):

Digital signal processing

OpenStudy (anonymous):

ohk

OpenStudy (anonymous):

So what step should I take next? In the problem

OpenStudy (anonymous):

W(Earth)/W(mars) = g(Earth)/g(mars) = [m(E)/m(M)] x [ R(E)/R(M)]^2 = 200/W

OpenStudy (anonymous):

now crossmultiply and evaluate W

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

let me check it out

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

make it[R(M)/R(E)]^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what does W represent in = 200/W

OpenStudy (anonymous):

is that the W I should isolate?

OpenStudy (anonymous):

Weight on mars

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

w=200/( [m(E)/m(M)] x [ R(E)/R(M)]^2 ) Am I right?

OpenStudy (anonymous):

R(m) /R(e)

OpenStudy (anonymous):

insteade of r(e) by r(m)

OpenStudy (anonymous):

you mean isolating and simplifying, I get w=r(e)/r(m)

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

no no....i mean replace the [R(e)/R(m)]^2 by [R(m)/R(e)]^2..rest is same

OpenStudy (anonymous):

in the first equaton?

OpenStudy (anonymous):

In this equation? W(Earth)/W(mars) = g(Earth)/g(mars) = [m(E)/m(M)] x [ R(E)/R(M)]^2 = 200/W

OpenStudy (anonymous):

W(Earth)/W(mars) = g(Earth)/g(mars) = [m(E)/m(M)] x [ R(M)/R(E)]^2 = 200/W

OpenStudy (anonymous):

from this equation I isolate W. and I should get the answer. Right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok thanks:)

OpenStudy (anonymous):

This is my email. emunrradtvamg@hotmail.com. If you have facebook or simething related, add me. Him, I have to leave now, I have to go home now. Im so angrey. Hope I can see you soon

OpenStudy (anonymous):

hungry, sorrry lol